OpenStudy (anonymous):

Enter the answer. 1. (1 pt) Which number replaces the box to make this a true sentence? 6 × (7 – 3) = (6 × 7) – ( × 3)

OpenStudy (texaschic101):

6 x (7 - 3) = (6 x 7) - (6 x 3) 6 (4) = 42 - 18 24 = 24 (correct)